OpenStudy (anonymous):

A price p (in dollars) and demand x for a product are related by 2x^2+1xp+50p^2=26000. If the price is increasing at a rate of 2 dollars per month when the price is 20 dollars, find the rate of change of the demand. Rate of change of demand =

OpenStudy (anonymous):

Step 1: Take the derivative of 2x^2+1xp+50p^2=26000 d/dt(2x^2+1xp+50p^2)=d/dt(26000)

OpenStudy (anonymous):

4xx'+x'p+xp'+100pp'=0 (4x+p)x'+(x+100p)p'=0 -(x+100p)p'/(4x+p)= x'

OpenStudy (anonymous):

Are you following?

OpenStudy (anonymous):

p=$20 2x^2+x(20)+50(20)^2=26000 (x+60)(x-50) so x= -60 and x =50 The value you select has to be greater than 0, so x=50

OpenStudy (anonymous):

yes yes i am following

OpenStudy (anonymous):

p=$20 p'=$2 per month x= 50 units Plug these values into x' x'=.........

OpenStudy (anonymous):

-((50+100(20))(2))/((4(50)+(20)))= x' x'=

OpenStudy (anonymous):

What do you get?

OpenStudy (anonymous):

i'll calculate

OpenStudy (anonymous):

is that negative((50.........?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

look at the original function

OpenStudy (anonymous):

x'= a negative

OpenStudy (anonymous):

is it -18.63?

OpenStudy (anonymous):

-18.64 actually?

OpenStudy (anonymous):

ya. -205/11 is the exact value

OpenStudy (anonymous):

YES!!! its correct

OpenStudy (anonymous):

thank you very very very much

OpenStudy (anonymous):

Welcome

OpenStudy (anonymous):

also another thankyou for explaining it